OpenStudy (anonymous):

A newly discovered planet orbits a distant star with the same mass as the Sun at an average distance of 130 million kilometers. Its orbital eccentricity is 0.4. How do you calculate the planet's orbital period?

OpenStudy (anonymous):

126,000,000km = ~.84AU P = R^(3/2) = .84^(3/2) = 0.769872717 P = .77 Years